### v1

@@ 4.1 Extract the solution space from DtN data
$T^{(t)}=T_{2}^{(t)}\circ T_{1}^{(t)},$
in Proposition 2.1, for any $\varphi\in\ker T_{1}^{(t)}$ we have
@@
This shows the inclusion
$\ker T_{1}^{(t)}\subseteq\ker T^{(t)}.$
Conversely, Proposition 3.1 guarantees that $T_{2}^{(t)}$ is injective. Therefore, for any $\varphi\in\ker T^{(t)}$, we have
@@ 4.2 Slice-by-slice reconstruction of conductivity
The base case begins with the layer that is closest to the origin. Note that $L_{d}=\{(1,1,\cdots,1)\}$ and $K_{d-1}=\{(1,\cdots,1,\underbrace{0}_{\text{the i th index}},1,\cdots,1):i=1,\cdots,d\}$. We begin the reconstruction of the conductivity $\gamma$ by focusing on the subset of edges connecting the node sets $K_{d-1}$ and $L_{d}$:
@@
Proof – Recovery of $\gamma$ on $E(K_{d-1},L_{d})$.
@@
Thus the subspace $\ker T_{1}^{(d-1)}$ uniquely determines the conductivity values $\gamma$ on the edge set $E(K_{d-1},L_{d})$. ∎
Next we decompose the edge set $E^{t}$ as
@@
collects all edges connecting nodes in $K_{i}\cup L_{i}$ to those in $K_{i+1}\cup L_{i+1}$. The interface graph
$G_{t}:=\bigl(L_{t}\cup J_{t}\cup L_{t+1},\ E_{t}\bigr)$ (4.3)
@@
Figure 4: $L_{t}$ (blue), $L_{t+1}$ (black), $K_{t}^{-}$ (orange) and $K_{t+1}^{+}$ (green) for $t\leq n+d-1$ (left) and t>n+d1t>n+d-1 (right) in the 3D case.
Inductive proof of Theorem 1.1: Suppose that the conductivity $\gamma$ on the edges $E^{t-1}$ is known (this holds true for the base case $t=d$). We impose the boundary potential by $\varphi\in\ker T_{1}^{(t)}$, and record the resulting boundary current $\Lambda_{\gamma}\varphi$. Given the conductivity on $E^{t-1}$, we can determine the interior potential distribution $\mathbf{u}$, which is supported within $L_{t}^{\mathcal{S}}$, where the existence of a unique solution is ensured by Lemma 3.1 . Then $\gamma|_{E_{t}}$ should satisfy ∑q∈N(p)γpq(up-uq)=0, for p∈Lt∪Lt+1,∑q∈N(p)γpq(up-uq)=ψp,  for p∈Jt. These relations can be expressed as a system of linear equations of the only unknown $\bm{\gamma}$:
$\mathbf{J}_{p}^{\mathbf{u}}\cdot\bm{\gamma}=C_{p},\text{ for all }p\in L_{t}\cup L_{t+1}\cup J_{t}{\text{ and }\mathbf{u}\in\mathcal{U}^{(t)},}$ (4.4)
where for each edge $e=pq$ in $E_{t}$, the vector $\mathbf{J}_{p}^{\mathbf{u}}\in\mathbb{R}^{E_{t}}$ is defined component-wise by
@@
The existence of a solution is always valid, and the solutions necessarily contain the true conductivity that generate the data, i.e., the DtN matrix $\Lambda_{\gamma}$. Furthermore, the next theorem gives the uniqueness of the solutions and thereby achieves the reconstruction of a single-layer conductivity.
@@
$\operatorname{span}\left(\left\{\mathbf{J}_{p}^{\mathbf{u}}:\mathbf{u}\in\mathcal{U}^{(t)},p\in L_{t}\cup L_{t+1}\cup J_{t}\right\}\right)^{\perp}=\{0\}.$
@@
We first investigate the space $\operatorname{span}\{\mathbf{J}_{p}^{\mathbf{u}}:\mathbf{u}\in\mathcal{U}^{(t)}\}$ for each node $p$ under three cases:
Figure 5: Illustration of $\mathbf{J}_{p}^{\mathbf{u}}$ with different $p$ in the 3D case with $L_{t}$ (blue), $L_{t+1}$ (black), $K_{t}^{-}$ (orange) and $K_{t+1}^{+}$ (green) .
Case 1: $p\in L_{t+1}$. Let $M_{p}=(L_{t}\cup K_{t})\cap\mathcal{N}(p)$, where $\mathcal{N}(p)$ is the set of neighbors of $p$. Consider the subspace
@@
If $\mathbf{w}\in\mathbb{R}^{M_{p}}$ is orthogonal to the subspace $\operatorname{span}\{\mathbf{u}|_{M_{p}}:\mathbf{u}\in\mathcal{U}^{(t)}\}$, then by Lemma 2.1, $(\mathbf{w},\mathbf{0})$ belongs to $\operatorname{span}\{\mathbf{v}_{q}|_{L_{t}}:q\in L_{t+1}^{\mathcal{S}}\}$ and supports in $M_{p}$. Corollary 3.2 implies $\mathbf{w}=\alpha\mathbf{v}_{p}|_{L_{t}}$ for some $\alpha\in\mathbb{R}$. By Lemma 3.4, the set $M_{p}$ contains exactly the nodes that are connected to $p$ in the graph $G_{t}$. Hence, by noting that $u_{p}=0$, we have
$\left(\operatorname{span}\{\mathbf{J}_{p}^{\mathbf{u}}:\mathbf{u}\in\mathcal{U}^{(t)}\}\right)^{\perp}=\operatorname{span}\{\bm{\gamma}|_{\mathcal{E}(p)}\}\times\mathbb{R}^{E\setminus\mathcal{E}(p)},$ (4.6)
where $\mathcal{E}(p)$ denotes edges incident to $p$, and $\bm{\gamma}|_{\mathcal{E}(p)}$ is supported on the edges directed toward $p$.
Case 2: $p\in L_{t}$ with no connection to $K_{t+1}^{+}$, or $p\in K_{t}^{-}$. Again, by Lemma 3.4, $\mathcal{N}(p)\cap G_{t}=\mathcal{N}(p)\cap L_{t+1}$. Furthermore, we have $\mathbf{u}|_{\mathcal{N}(p)\cap L_{t+1}}=0$. Let $\mathbf{1}_{\mathcal{E}(p)}$ denote the indicator vector on edges incident to $p$. Therefore, to prove
@@
it suffices to show that there exists $\mathbf{u}\in\mathcal{U}^{(t)}$ with $\mathbf{u}_{p}\neq 0$. Indeed, if no such $\mathbf{u}$ exists, then the canonical basis vector $\mathbf{e}_{p}\in\mathbb{R}^{J_{t}^{\mathcal{S}}\cup L_{t}^{\mathcal{S}}}$ is orthogonal to $\mathcal{U}^{(t)}$ and supported on some $M_{q}$ with $q\in\mathcal{N}(p)\cap L_{t+1}$. By Corollary 3.2, this implies $\mathbf{e}_{p}=\alpha\mathbf{v}_{q}|_{L_{t}}$ for some $\alpha\in\mathbb{R}$, which contradicts the positivity of $\gamma$.
Case 3: $p\in L_{t}$ connected to $q\in K_{t+1}^{+}$. By an argument analogous to Case 2, one readily verifies that
$\operatorname{span}\{\mathbf{J}_{p}^{\mathbf{u}}:\mathbf{u}\in\mathcal{U}^{(t)}\}+\operatorname{span}\{\mathbf{J}_{q}^{\mathbf{u}}:\mathbf{u}\in\mathcal{U}^{(t)}\}=\underbrace{\mathbb{R}}_{\text{edge }pq}\times\operatorname{span}\{\mathbf{1}_{\mathcal{E}(p)\setminus pq}\}\times\underbrace{\mathbf{0}}_{\text{edges }E\backslash\mathcal{E}(p)},$ (4.8)
where $\mathbf{1}_{\mathcal{E}(p)\setminus pq}$ denotes the indicator vector on edges incident to $p$ excluding the edge $pq$.
Now suppose that $\mathbf{w}$ lies in the orthogonal complement, i.e.,
@@
We claim $\mathbf{w}=\mathbf{0}$. First, fix any node $p\in L_{t+1}$. Since $\mathbf{w}$ is orthogonal to each $\mathbf{J}_{p}^{\mathbf{u}}$, its restriction to edges incident to $p$ is proportional to $\bm{\gamma}|_{\mathcal{E}(p)}$ as in (4.6). Hence, there exists $c_{p}\in\mathbb{R}$ such that
@@
Next, the orthogonality condition $\mathbf{w}\perp\mathbf{J}_{p}^{\mathbf{u}}$ for all $p\in L_{t}\cup J_{t}$, together with (4.8) and (4.7), implies the linear constraints
@@
We extend $\mathbf{c}$ by setting $\mathbf{c}=0$ on $L_{t}^{\mathcal{S}}\cup J_{t}^{\mathcal{S}}$, so that the boundary nodes $J_{t}^{\mathcal{S}}$ have zero potential and zero current. Then $\mathbf{c}$ solves the boundary value problem
$\left\{\begin{aligned} \Delta_{\gamma}\mathbf{c}&=\mathbf{0}\quad\text{in }L_{t},\\ \mathbf{c}&=0\quad\text{on }J_{t}^{\mathcal{S}},\\ D_{\gamma}\mathbf{c}&=0\quad\text{on }J_{t}^{\mathcal{S}}.\end{aligned}\right$.
By Lemma 3.1, the only solution is $\mathbf{c}=0$ on $L_{t+1}$. This and the identity (4.10) imply that $\mathbf{w}$ vanishes on all edges incident to nodes in $L_{t+1}$, except possibly those incident to nodes in $K_{t+1}^{+}$. Finally, by (4.8), we conclude that $\mathbf{w}=\mathbf{0}$ everywhere. This completes the proof of the theorem. ∎
@@
The data pairs
@@
uniquely determine the conductivity $\bm{\gamma}$ on the entire edge set $E$.
@@
The true conductivity $\bm{\gamma}$ lies in the set {γJ_p^uγ=C_p for all p∈L_t∪L_t+1∪J_t,uU^(t)}. If $\bm{\gamma}_{1}$ and $\bm{\gamma}_{2}$ lie in the set, by Theorem 4.1, we have
$\gamma_{1}-\gamma_{2}\in\operatorname{span}\left(\left\{\mathbf{J}_{p}^{\mathbf{u}}:\mathbf{u}\in\mathcal{U}^{(t)},p\in L_{t}\cup L_{t+1}\cup J_{t}\right\}\right)^{\perp}=\{0\},$
which establishes uniqueness. ∎
Upon combining Lemma 4.1 and Theorem 4.2, the proof of Theorem 1.1 is complete. $\blacksquare$

### v2

@@ 4.1 Extract the solution space from the DtN data $\Lambda_{\gamma}$
$T^{(t)}=T_{2}^{(t)}\circ T_{1}^{(t)}$
from Proposition 2.1, for any $\varphi\in\ker T_{1}^{(t)}$, we have
@@
This shows the inclusion $\ker T_{1}^{(t)}\subseteq\ker T^{(t)}$. Conversely, Proposition 3.1 guarantees that $T_{2}^{(t)}$ is injective. Therefore, for any $\varphi\in\ker T^{(t)}$, we have
@@ 4.2 Slice-by-slice reconstruction of conductivity
In this part we present the proof of Theorem 1.1. It is based on mathematical induction along the slices. The base case begins with the layer that is closest to the origin. Note that $L_{d}=\{(1,1,\cdots,1)\}$ and $K_{d-1}=\{(1,\cdots,1,\underbrace{0}_{\text{the $i$th index}},1,\cdots,1):i=1,\cdots,d\}$. We begin the reconstruction of the conductivity $\gamma$ by focusing on the subset of edges connecting the node sets $K_{d-1}$ and $L_{d}$:
@@
Lemma 4.2.
The DtN matrix $\Lambda_{\gamma}$ uniquely determines $\gamma$ on $E(K_{d-1},L_{d})$.
Proof.
@@
Thus the subspace $\ker T_{1}^{(d-1)}$ uniquely determines the conductivity $\gamma$ on the edge set $E(K_{d-1},L_{d})$. ∎
Next we define an edge set $E^{t}$ by
@@
collects all edges connecting nodes in $K_{i}\cup L_{i}$ to those in $K_{i+1}\cup L_{i+1}$. With $\emptyset$ being the empty set, the interface graph
$G_{t}:=\bigl(E_{t},\emptyset,L_{t}\cup J_{t}\cup L_{t+1}\bigr)$ (4.3)
@@
Figure 4: $L_{t}$ ((blue)), $L_{t+1}$ ((black)), $K_{t}^{-}$ ((orange)) and $K_{t+1}^{+}$ ((green)) for $t\leq n+d-1$ ((left)) and t>n+d1t>n+d-1 ((right)) in the 3D case.
Inductive proof of Theorem 1.1: Suppose that the conductivity $\gamma$ on the edges $E^{t-1}$ is known (this holds true for the base case $t=d$, cf. Lemma 4.2). We impose the boundary potential by $\varphi\in\ker T_{1}^{(t)}$, and record the resulting boundary current $\psi=\Lambda_{\gamma}\varphi$. Given the conductivity $\gamma$ on $E^{t-1}$, by Lemma 3.1, we can determine the interior potential distribution $\mathbf{u}$, which is supported within $L_{t}^{\mathcal{S}}$. Then the subvector $\widetilde{\gamma}:=\gamma|_{E_{t}}$ should satisfy ∑q∈N(p)γpq(up-uq)=0, for p∈Lt∪Lt+1,∑q∈N(p)γpq(up-uq)=ψp,  for p∈Jt. These relations can be written into a system of linear equations of the only unknown $\widetilde{\gamma}$:
$\mathbf{J}_{p}^{\mathbf{u}}\cdot\widetilde{\gamma}=C_{p},\quad\forall p\in L_{t}\cup L_{t+1}\cup J_{t}{\text{ and }\mathbf{u}\in\mathcal{U}^{(t)},}$ (4.4)
where $C_{p}=\psi_{p}$ for $p\in J_{t}$ and $C_{p}=0$ otherwise, and for each edge $e=pq$ in $E_{t}$, the vector $\mathbf{J}_{p}^{\mathbf{u}}\in\mathbb{R}^{E_{t}}$ is defined componentwise by (see Fig. 5)
@@
The existence of a solution to the system (4.4) is always ensured (by the assumption that it is realized by the true conductivity), and the solutions necessarily contain the true conductivity that generates the DtN matrix $\Lambda_{\gamma}$. Furthermore, the next theorem gives the uniqueness of the solutions and thereby gives the reconstruction of a single-layer conductivity.
@@
$\operatorname{span}\left(\left\{\mathbf{J}_{p}^{\mathbf{u}}:\mathbf{u}\in\mathcal{U}^{(t)},p\in L_{t}\cup L_{t+1}\cup J_{t}\right\}\right)^{\perp}=\{\mathbf{0}\}.$ (4.6)
@@
We first investigate the space $\operatorname{span}\{\mathbf{J}_{p}^{\mathbf{u}}:\mathbf{u}\in\mathcal{U}^{(t)}\}$ for each node $p\in L_{t}\cup L_{t+1}\cup J_{t}$ under the following three cases: (i) $p\in L_{t+1}$, (ii) $p\in L_{t}$ with no connection to $K_{t+1}^{+}$, or $p\in K_{t}^{-}$, and (iii) $p\in L_{t}$ connected to $q\in K_{t+1}^{+}$. Below we analyze the three cases separately.
Figure 5: A schematic illustration of $\mathbf{J}_{p}^{\mathbf{u}}$ with different $p$ in the 3D case with $L_{t}$ (blue), $L_{t+1}$ ((black)), $K_{t}^{-}$ ((orange)) and $K_{t+1}^{+}$ ((green)).
Case (i): $p\in L_{t+1}$. Let $M_{p}=(L_{t}\cup J_{t})\cap\mathcal{N}(p)$, where the set $\mathcal{N}(p)$ consists of the neighbors of $p$. Consider the subspace
@@
If $\mathbf{w}\in\mathbb{R}^{M_{p}}$ is orthogonal to the subspace $\operatorname{span}\{\mathbf{u}|_{M_{p}}:\mathbf{u}\in\mathcal{U}^{(t)}\}$, then by Lemma 2.1, the zero extension, still denoted by $\mathbf{w}\in\mathbb{R}^{L_{t}^{\mathcal{S}}\cup J_{t}^{\mathcal{S}}}$, belongs to $\operatorname{span}\{\mathbf{v}_{q}|_{L_{t}^{\mathcal{S}}\cup J_{t}^{\mathcal{S}}}:q\in L_{t+1}^{\mathcal{S}}\}$ and is supported in $M_{p}$. Corollary 3.2 implies $\mathbf{w}=\alpha\mathbf{v}_{p}|_{L_{t}^{\mathcal{S}}\cup J_{t}^{\mathcal{S}}}$ for some $\alpha\in\mathbb{R}$. Meanwhile, by noting $u_{p}=0$, we have
$\left(\operatorname{span}\{\mathbf{J}_{p}^{\mathbf{u}}:\mathbf{u}\in\mathcal{U}^{(t)}\}\right)^{\perp}=\operatorname{span}\{\bm{\gamma}|_{\mathcal{E}(p)}\}\times\mathbb{R}^{E_{t}\setminus\mathcal{E}(p)},$ (4.7)
where $\mathcal{E}(p):=E(p,M_{p})$ denotes the set of edges incident to $p$.
Case (ii): $p\in L_{t}$ with no connection to $K_{t+1}^{+}$, or $p\in K_{t}^{-}$. By Lemma 3.4, $\mathcal{N}(p)\cap(L_{t}\cup J_{t}\cup L_{t+1})=\mathcal{N}(p)\cap L_{t+1}$. Furthermore, we have $\mathbf{u}|_{\mathcal{N}(p)\cap L_{t+1}}=0$ for any $\mathbf{u}\in\mathcal{U}^{(t)}$. Let $\mathbf{1}_{\mathcal{E}(p)}$ denote the indicator vector of $\mathcal{E}(p)$ (i.e., edges incident to $p$). Then to prove
@@
it suffices to show that there exists $\mathbf{u}\in\mathcal{U}^{(t)}$ with $\mathbf{u}_{p}\neq 0$. Indeed, if no such $\mathbf{u}$ exists, then the canonical basis vector $\mathbf{e}_{p}\in\mathbb{R}^{J_{t}^{\mathcal{S}}\cup L_{t}^{\mathcal{S}}}$ is orthogonal to $\mathcal{U}^{(t)}$ and naturally supported on some $M_{q}$ with $q\in\mathcal{N}(p)\cap L_{t+1}$. By Corollary 3.2, this implies $\mathbf{e}_{p}=\alpha\mathbf{v}_{q}|_{J_{t}^{\mathcal{S}}\cup L_{t}^{\mathcal{S}}}$ for some $\alpha\in\mathbb{R}$. However, for any positive $\gamma$, the vector $\mathbf{v}_{q}|_{J_{t}^{\mathcal{S}}\cup L_{t}^{\mathcal{S}}}$ contains more than one nonzero entries, whereas $\mathbf{e}_{p}$ contains only one nonzero entry, leading to a contradiction.
Case (iii): $p\in L_{t}$ connected to $q\in K_{t+1}^{+}$. By an argument analogous to Case (ii), one can readily verify that
$\operatorname{span}\{\mathbf{J}_{p}^{\mathbf{u}}:\mathbf{u}\in\mathcal{U}^{(t)}\}+\operatorname{span}\{\mathbf{J}_{q}^{\mathbf{u}}:\mathbf{u}\in\mathcal{U}^{(t)}\}$ (4.9)
$=$ $\underbrace{\mathbb{R}}_{\text{edge }pq}\times\operatorname{span}\{\mathbf{1}_{\mathcal{E}(p)\setminus\{pq\}}\}\times\underbrace{\mathbf{0}}_{\text{edges }E_{t}\backslash\mathcal{E}(p)},$
where the notation $\mathbf{1}_{\mathcal{E}(p)\setminus\{pq\}}$ denotes the indicator vector on edges incident to $p$ excluding the edge $pq$.
Now we can prove the desired assertion (4.6). Suppose that $\mathbf{w}$ lies in the orthogonal complement, i.e.,
@@
We claim $\mathbf{w}=\mathbf{0}$. Since for any node $p\in L_{t+1}$, $\mathbf{w}$ is orthogonal to each $\mathbf{J}_{p}^{\mathbf{u}}$, its restriction $\mathbf{w}|_{\mathcal{E}(p)}$ is proportional to $\bm{\gamma}|_{\mathcal{E}(p)}$ as in (4.7). Hence for any node $p\in L_{t+1}$, there exists $c_{p}\in\mathbb{R}$ such that
@@
Next, the orthogonality condition $\mathbf{w}\perp\mathbf{J}_{p}^{\mathbf{u}}$ for all $p\in L_{t}\cup J_{t}$, together with the relations (4.8) and (4.9), leads to the following system of linear constraints
@@
We extend the vector $\mathbf{c}$ by setting $\mathbf{c}=0$ on $L_{t}^{\mathcal{S}}\cup J_{t}^{\mathcal{S}}$, so that the boundary nodes $J_{t}^{\mathcal{S}}$ have zero potential and zero current. Then $\mathbf{c}$ solves the boundary value problem
$\left\{\begin{aligned} \Delta_{\gamma}\mathbf{c}&=\mathbf{0}\quad\text{in }L_{t}^{\mathcal{S}},\\ \mathbf{c}&=\mathbf{0}\quad\text{on }J_{t}^{\mathcal{S}},\\ D_{\gamma}\mathbf{c}&=\mathbf{0}\quad\text{on }J_{t}^{\mathcal{S}}.\end{aligned}\right$.
By Lemma 3.1, the only solution is $\mathbf{c}=0$ on $L_{t+1}^{\mathcal{S}}$. This and the identity (4.11) imply that $\mathbf{w}$ vanishes on all edges incident to nodes in $L_{t+1}$, except possibly those incident to nodes in $K_{t+1}^{+}$. However, by (4.9), we can finally conclude that $\mathbf{w}=\mathbf{0}$ everywhere. This completes the proof of the theorem. ∎
@@
Given the conductivity $\gamma$ on the edge set $E^{t-1}$, the data pairs
@@
uniquely determine the conductivity $\gamma$ on the edge set $E_{t}$.
@@
The true conductivity $\bm{\gamma}:=\gamma|_{E_{t}}$ lies in the set
$\{\bm{\gamma}\mid\mathbf{J}_{p}^{\mathbf{u}}\cdot\bm{\gamma}=C_{p}\text{ for all }p\in L_{t}\cup L_{t+1}\cup J_{t},\mathbf{u}\in\mathcal{U}^{(t)}\},$
with $C_{p}$ from (4.4). If $\bm{\gamma}_{1}$ and $\bm{\gamma}_{2}$ both lie in the set, by Theorem 4.1, we have
$(\gamma_{1}-\gamma_{2})|_{E_{t}}\in\operatorname{span}\left(\left\{\mathbf{J}_{p}^{\mathbf{u}}:\mathbf{u}\in\mathcal{U}^{(t)},p\in L_{t}\cup L_{t+1}\cup J_{t}\right\}\right)^{\perp}=\{\mathbf{0}\},$
which establishes the desired uniqueness. ∎
Combining Lemma 4.1 and Theorem 4.2 completes the proof of Theorem 1.1. $\blacksquare$
Remark 4.1.
For two-dimensional square lattices, Curtis and Morrow [20] provided a sharp characterization on a subset of entries of the DtN matrix $\Lambda_{\gamma}$ that is necessary and sufficient for the recovery of the full DtN matrix $\Lambda_{\gamma}$, and using this important fact, established the unique determination of the conductivity $\gamma$ by a partial DtN matrix. For multi-dimensional hypercubic lattices, in view of the symmetry of the lattice structure, one naturally expects the existence of a subset of entries of the DtN matrix $\Lambda_{\gamma}$ for the unique recovery of $\Lambda_{\gamma}$. However, a precise characterization of the minimal subset of $\Lambda_{\gamma}$ that is necessary and sufficient for the unique recovery is still missing. Note also that the present analysis does not extend directly to the case of partial DtN matrices.
Remark 4.2.
The discrete Calderón problem on cylindrical networks studied by Lam and Pylyavskyy [25] exhibits inherent non-uniqueness using the invariance of the DtN matrix under the so-called $Y$-$\Delta$ transformation. The key distinction between the cylindrical setting and hyper-cubic lattices is topological: cylindrical networks reside on a cylinder (or torus), whereas hyper-cubic lattices live on bounded domains in Euclidean spaces. In the Euclidean case, bounded regions possess more boundary nodes than the torus, which leads to different uniqueness results. Cylindrical networks can be reduced to networks in bounded domains in Euclidean spaces by cutting the cylinder along altitude, which gives rise to two families of boundary points along the cuts. The extension of the cylindrical setting to higher dimensions, in which networks are embedded on a higher-dimensional torus, presents intriguing possibilities that deserve further investigations.